OpenStudy (anonymous):

2x-6y=3 X+4y=5 By substitution

OpenStudy (anonymous):

Need help please

OpenStudy (anonymous):

x=5-4y Plug this into the other equation 2(5-4y)-6y=3 10-8y-6y=3 -14y=-7 y=1/2 x=5-4(1/2) x=3

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

No problem!